given A(18,9), B(6,27), and C(6,9), determine the coordinates of point D such that CD is parallel to AB and D is on the:
i)y-axis
ii)x-axis

Answer by richwmiller(17219)

A(18,9), B(6,27), and C(6,9)
D(0,y) on the y axis
D(x,0) on the x axis
y = 36-(3x)/2
y = 36-(3x)/2
9 = b-(3*6)/2
9 = b-9
b=18
y = 18-(3x)/2
D(0,18) on the y axis
y = 18-(3x)/2
y=0
3x/2=18
3x=36
x=12
D(12,0) on the x axis